# Normalized mutual information is a biased measure for classification and community detection

### v2

@@ I Introduction
A common task in data analysis is the comparison of two different labelings of the same set of objects. How well do demographics predict political affiliation? How accurately do blood tests predict clinical outcomes? The information theoretic measure known as mutual information is commonly applied to answer questions like these, in which an experimental or computational estimate of some kind is compared against a “ground truth.”
@@
Mutual information [1] is a measure of how efficiently we can describe one labeling of the set of objects if we already know another labeling. Specifically, it measures how much less information it takes to communicate the first labeling if we know the second versus if we do not. As an example, mutual information is commonly used in network science to evaluate the performance of algorithms for network community detection [2]. One takes a network whose community structure is already known and applies a community detection algorithm to it to infer the communities. Then one uses mutual information to compare the output of the algorithm to the known correct communities. Algorithms that consistently achieve high mutual information scores are considered good. We will use this application as an illustrative example later in the paper.
@@
Mutual information has a number of appealing properties as a tool for comparing labelings. It is invariant under permutations of the labels, so that, for example, two divisions of a network into communities that differ only in how the communities are labeled will be correctly identified as being the same. It also returns sensible results in cases where the number of distinct label values is not the same in the two labelings. However, the standard mutual information measure also has some significant shortcomings, and two in particular that we highlight in this paper. First, it has a bias towards labelings with too many distinct label values. For instance, a community detection algorithm that routinely divides networks into significantly more groups than are present in the ground truth can nonetheless achieve high mutual information scores. A number of approaches for correcting this flaw have been proposed. One can apply direct penalties for incorrect numbers of groups [3] or subtract correction terms based on the average value of the mutual information over some ensemble of candidate labelings [4, 5] or on the statistics of the contingency table [6, 7]. For reasons we discuss shortly, we favor the latter approach, which leads to the measure known as the reduced mutual information.
@@
The second drawback of the mutual information arises when the measure is normalized, as it commonly is to improve interpretability. The most popular normalization scheme creates a measure that runs between zero and one by dividing the mutual information by the arithmetic mean of the entropies of the two labelings being compared [8], although one can also normalize by the minimum, maximum, or geometric mean of the entropies. As we demonstrate in this paper, however, these normalizations introduce biases into the results by comparison with the unnormalized measure, because the normalization factor depends on the candidate labeling as well as the ground truth. This effect can be large enough to change scientific conclusions, and we provide examples of this phenomenon.
@@
In order to avoid this latter bias, while still retaining the interpretability of a normalized mutual information measure, in this paper we propose normalizing by the entropy of the ground-truth labeling alone. This removes the source of bias but introduces an asymmetry in the normalization. At first sight this asymmetry might seem undesirable, and previous authors have gone to some lengths to avoid it. Here, however, we argue that it is not only justified but even desirable, for several reasons. First, many of the classification problems we consider are unaffected by the asymmetry, since they involve the comparison of one or more candidate labelings against a single, unchanging ground truth. Moreover, by contrast with the multitude of possible symmetric normalizations, the asymmetric measure we propose is the unique unbiased way to normalize the mutual information such that a perfect match receives a score of 1.
@@
But more broadly, one can argue that an asymmetric measure can be more informative than the conventional symmetric one. Consider for instance the common situation where one labeling is a fine-grained version of the other. We might for example label individuals by the country they live in on the one hand and by the town or city on the other. Then the detailed labeling tells us everything there is to know about the coarse-grained one, but the reverse is not true: telling you the city fixes the country, but not vice versa. Thus one could argue that the mutual information between the two should be asymmetric even before normalization—labeling A contains more information about labeling B than B does about A—and this type of asymmetry will be a key feature of the measures we study.
@@
Both drawbacks of the standard mutual information—bias towards too many groups and dependence of the normalization on the candidate labeling—can be addressed simultaneously by using an asymmetric normalized reduced mutual information as defined in this paper. In support of this approach we present an extensive comparison of the performance of this and other variants of the mutual information in network community detection tasks, generating a large number of random test networks with known community structure and a variety of structural parameters, and then attempting to recover the communities using popular community detection algorithms. Within this framework, we find that conclusions about which algorithms perform best are significantly impacted by the choice of mutual information measure, and specifically that traditional measures erroneously favor algorithms that find too many communities, but our proposed measure does not. Code implementing our measure may be found at https://github.com/maxjerdee/reduced_mutual_information.
II Mutual information
Mutual information can be thought of in terms of the amount of information it takes to transmit a labeling from one person to another. In general, if we have a set of $N$ possible values for a quantity, such as possible labelings of a set of objects, then one can communicate a single one of them to a receiver by assigning each value a unique binary string and then transmitting the appropriate string. The minimum number of bits needed to do this is
$H=\lceil\log_{2}N\rceil\simeq\log_{2}N\propto\ln N.$ (1)
Conventionally one uses base-2 logarithms in this context, which gives $H$ units of bits, but we use natural logarithms in this paper, since they are more convenient for our purposes. The only difference is an overall factor of $\ln 2$, which will have no effect on our conclusions. Henceforth, all logarithms should be read as natural logs.
@@ II Mutual information
Suppose that we have a labeling or division of $n$ objects into $q$ groups, which we represent as a vector of $n$ integer elements, each with value in the range $1\ldots q$. We assume there to be a ground-truth labeling, which we denote $g$, and a candidate labeling for comparison $c$, generated for instance by some sort of algorithm. The mutual information $I(c;g)$ between the two is the amount of information that can be saved when transmitting the truth $g$ if the receiver already knows the candidate $c$. We can write this information as the total entropy of $g$ minus the conditional entropy of $g$ given $c$:
$I(c;g)=H(g)-H(g|c).$ (2)
@@
Loosely, we say that $I(c;g)$ measures “how much $c$ tells us about $g$.”
We now write down expressions for both entropies in (2), but there are some subtleties involved in doing so, because the expressions depend on how the transmitted information is encoded and also because achieving correct results requires us to carefully retain some terms that are often neglected. Our presentation follows that of [7]. First consider $H(g)$, which represents the information required to transmit the ground truth on its own. The transmission process has three steps. We first transmit the number of groups $q_{g}$ in the labeling, followed by a vector $n^{(g)}$ with $q_{g}$ elements $n^{(g)}_{s}$ equal to the number of objects in each group. Third, we transmit our final labeling $g$. If we apply Eq. (1) to calculate the cost of the first step, since the maximum possible value of $q_{g}$ is $n$, transmitting any particular value requires information $H(q_{g})=\log n$. Likewise for the third step, given $n^{(g)}$, there are $n!/\prod_{s}n_{s}^{(g)}!$ labelings $g$ that have the correct group sizes, so the amount of information needed to identify a particular labeling uniquely is
$H(g|n^{(g)})=\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}.$ (3)
The middle step, transmission of the group sizes, is more delicate: here some optimization is possible that allows us to achieve a better measure of information content. Generally when transmitting a sequence of $n$ values with unequal probabilities such that value $r$ occurs $n_{r}$ times, the information cost is given by the generalization of Eq. (3) thus:
$\log{n!\over\prod_{r}n_{r}!}$ $\simeq n\log n-n-\sum_{r}(n_{r}\log n_{r}-n_{r})$
$=-\sum_{r}n_{r}\log p_{r},$ (4)
where $p_{r}=n_{r}/n$ is the probability of value $r$ and we have approximated the factorials using Stirling’s formula. Equation 4 tells us that the information cost to transmit the value $r$ is simply $-\log p_{r}$. If the $N$ possible values are equally likely to appear and $p_{r}=1/N$, we recover the information cost of Eq. (1), but if they are not we can take advantage of that fact to create more efficient encodings.
The transmission of $n^{(g)}$ benefits from such an efficient encoding. We do not in general know the distribution of the elements of $n^{(g)}$, but we know they that may be—and often are—heterogeneous. As shown in [7], an effective approach is to assume a Dirichlet-multinomial distribution with concentration parameter $\alpha_{g}>0$
$P(n^{(g)}|q_{g},\alpha_{g})=$ (5)
$\qquad\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}^{-1}\prod_{r=1}^{q_{g}}\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1},$
where for non-integer $\alpha_{g}$ we generalize the binomial coefficient in the obvious way:
$\Bigl{(}{n\atop k}\Bigr{)}={\frac{\Gamma(n+1)}{\Gamma(k+1)\Gamma(n-k+1)}},$ (6)
with $\Gamma(x)$ being the standard gamma function. If $\alpha_{g}=1$, the Dirichlet-multinomial distribution is uniform over all $q_{g}$-vectors $n^{(g)}$ of non-negative integers that sum to $n$. If $10<\alpha_{g}<$ 1 the distribution favors vectors with more extreme elements, while for $\alpha_{g}>1$ it favors vectors with more homogeneous elements. With this choice the resulting information cost to transmit $n^{(g)}$ is
$H($ $n^{(g)}|q_{g},\alpha_{g})=-\log P(n^{(g)}|q_{g},\alpha_{g})$ (7)
$=\log\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}-\sum_{r=1}^{q_{g}}\log\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1}.$
The optimal encoding for transmitting $n^{(g)}$ within the Dirichlet-multinomial family is given by the minimum of this expression with respect to $\alpha_{g}$, which is also equivalent to simply maximizing $P(n^{(g)}|\alpha_{g})$, i.e., to finding the maximum-likelihood value of $\alpha_{g}$. Note that we omit the cost of transmitting the parameter $\alpha_{g}$ itself since it will ultimately cancel out of the expression for the mutual information as described in [7].
The total cost to transmit $g$ is now equal to the sum of the individual costs of the three steps, thus:
$H(g)$ $=H(q_{g})+H(n^{(g)}|q_{g},\alpha_{g})+H(g|n^{(g)})$
$=\log n+\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}$ (8)
$+\log\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}-\sum_{r=1}^{q_{g}}\log\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1}$
This three-part encoding scheme is not the only one that can be applied to this problem, but it is an efficient one in the common case of a relatively small number of groups $q_{g}\ll n$ with potentially unequal sizes, and it reduces to the conventional expression for $H(g)$ when subleading terms are neglected. It is an important part of our argument here, however, that the subleading terms be retained and carried through to the calculation of the mutual information.
For the calculation of $H(g|c)$, the second term in Eq. (2), we use a similar three-part encoding, but one that takes advantage of the receiver’s knowledge of $c$ when communicating $g$. We first communicate $q_{g}$ as before, at the same information cost of $H(q_{g})=\log n$. In place of the vector $n^{(g)}$ we then communicate a contingency table $n^{(gc)}$, a matrix with elements $n_{rs}^{(gc)}$ equal to the number of objects that simultaneously belong to group $r$ in the candidate labeling $c$ and group $s$ in the ground truth $g$. Figure 1 shows an example of a contingency table for two labelings of the nodes in a small network.
Figure 1: An example of a contingency table for two colorings of the same network, one with three colors and one with two. The entries in the $3\times 2$ contingency table $n^{(gc)}$ count the number of nodes that have each combination of colors. The row and column sums $n^{(g)}$ and $n^{(c)}$ of the contingency table count the number of nodes with each color in the two colorings. Note that, although we illustrate the contingency table with an application to a network, the table itself is independent of the network structure.
Our process for transmitting the contingency table is analogous to the one we use for transmitting $n^{(g)}$. Since the labeling $c$ is known, the column sums $n^{(c)}$ of the contingency table are known as well, and we assume each column to be drawn from a Dirichlet-multinomial distribution with the appropriate sum. We use the same concentration parameter $\alpha_{g|c}$ for every column, but the columns are otherwise independent. If we denote column $s$ by $n_{\cdot s}^{(gc)}$, then the information cost to transmit the entire contingency table is
$H($ $n^{(gc)}|n^{(c)},q_{g},\alpha_{g|c})=\sum_{s=1}^{q_{c}}H(n_{\cdot s}^{(gc)}|n_{s}^{(c)},q_{g},\alpha_{g|c})$
$=\sum_{s=1}^{q_{c}}\biggl{[}\log\binom{n_{s}^{(c)}+q_{g}\alpha_{g|c}-1}{q_{g}\alpha_{g|c}-1}$
$\hskip 60.00009pt-\sum_{r=1}^{q_{g}}\log\binom{n_{rs}^{(gc)}+\alpha_{g|c}-1}{\alpha_{g|c}-1}\biggr{]}.$ (9)
This formulation allows for very efficient transmission of the contingency table under commonly occurring conditions. Consider for instance the special (but not implausible) case where $g=c$, so that $n^{(gc)}$ is diagonal. In this case the optimal concentration parameter is $\alpha_{g|c}=0$, which restricts the possible contingency tables to those for which each column $n_{\cdot s}^{(gc)}$ has exactly one nonzero entry. Given that we know the column sums, this leaves only the positions of the nonzero entries within each column to be transmitted, which results in near-optimal performance, a decisive improvement over previous reductions of the mutual information [7]. Most practical applications will not have $c=g$ precisely, but the two labelings are often similar and similarly impressive improvements in performance can be realized in such cases as well.
Finally, having transmitted the contingency table, it remains only to transmit the ground-truth labeling itself, where we need consider only those labelings consistent with the contingency table, given the known candidate labeling $c$. The number of such labelings is $\prod_{r}n_{r}^{(c)}!/\prod_{rs}n_{rs}^{(gc)}!$, so the information needed to uniquely identify one of them is
$H(g|n^{(gc)},c)=\log\frac{\prod_{r}n_{r}^{(c)}!}{\prod_{rs}n_{rs}^{(gc)}!}.$ (10)
In typical applications the number of labelings compatible with the contingency table is much smaller than the total number of labelings, and hence this encoding substantially reduces the amount of information needed to transmit the ground truth. It is not the only encoding possible, but it is an efficient one in practice and it is the one upon which the conventional mutual information is based.
Collecting all our terms, the total conditional information can now be written
$H(g|c)$ $=H(q_{g})+H(n^{(gc)}|n^{(c)},q_{g},\alpha_{g|c})+H(g|n^{(cg)},c).$ (11)
And substituting (8) and (11) into Eq. (2), we get a complete expression for the mutual information
@@
$=\log\frac{n!\prod_{rs}n_{rs}^{(gc)}!}{\prod_{r}n_{r}^{(c)}!\prod_{s}n_{s}^{(g)}!}+\log\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}$
$\qquad-\sum_{r=1}^{q_{g}}\log\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1}$
$\qquad-\sum_{s=1}^{q_{c}}\log\binom{n_{s}^{(c)}+q_{g}\alpha_{g|c}-1}{q_{g}\alpha_{g|c}-1}$
$\qquad+\sum_{r=1}^{q_{g}}\sum_{s=1}^{q_{c}}\log\binom{n_{rs}^{(gc)}+\alpha_{g|c}-1}{\alpha_{g|c}+1}.$ (12)
In most treatments, only the first of these terms is retained, since the others are subdominant and often small [6, 7]. When all the terms are retained, however, the quantity (12) is known as the reduced mutual information and there are good reasons to use this full expression in certain circumstances, a point that we investigate in Section II.1.
If we follow convention, however, and keep only the leading term, then we arrive at the traditional mutual information, which we denote $I_{0}$:
@@
For later convenience, we also define entropies $H_{0}$ for the individual labelings ignoring subdominant terms, which can be thought of as the mutual information between the labelings and themselves:
$H_{0}(c)$ $=I_{0}(c;c)=\log\frac{n!}{\prod_{r}n_{r}^{(c)}!},$ (14)
$H_{0}(g)$ $=I_{0}(g;g)=\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}.$ (15)
Note that these differ from the entropy of Eq. (8), which includes subdominant terms. Trivially (14) and (15) provide upper bounds on the mutual information, while zero is a lower bound, so
$0\leq I_{0}(c;g)\leq H_{0}(c),H_{0}(g).$ (16)
These bounds will be useful shortly.
A common further step is to approximate the mutual information of Eq. (13) using Stirling’s approximation in the form
$\log n!=n\log n-n+\mathrm{O}(\log(n)),$ (17)
which gives
$I_{0}(c;g)=n\sum_{rs}p_{rs}^{(gc)}\log\frac{p_{rs}^{(gc)}}{p_{r}^{(c)}p_{s}^{(g)}}+\mathrm{O}\left(\log n\right),$ (18)
where
$p_{rs}^{(gc)}={n_{rs}^{(gc)}\over n},\qquad p_{r}^{(c)}={n_{r}^{(c)}\over n},\qquad p_{s}^{(g)}={n_{s}^{(g)}\over n},$ (19)
and we adopt the convention that $00\log 0=$ 0. Equation (18) is often used as the definition of the mutual information, although technically it is only an approximation, albeit one that is asymptotically correct in the limit of large $n$. (It can also be derived by considering the information cost to transmit the true label of a single object, selected at random, with and without knowledge of the corresponding candidate label.)
II.1 Reduced mutual information
Equation (13) defines the standard mutual information. In this paper we explore variant definitions that can perform better at the clustering tasks we consider. A number of variants have been proposed, of which the simplest in the context of the current presentation is the reduced mutual information $I(c;g)$ of Eq. (12), which is simply the value of the mutual information when one does not neglect the subleading information cost of transmitting the group sizes and the contingency table.
To understand the importance of the subleading terms let us explore how things can go wrong if we neglect them. To take an extreme example, suppose an algorithm simply places every object in a group of its own, which corresponds to the candidate labeling $c=(1,\ldots,n)$. No matter what the ground truth is, this choice of $c$ clearly contains no information about it whatsoever, so we might expect the mutual information to be zero. This, however, is not what we find. The contingency table in this case is
$n_{rs}^{(gc)}=\biggl{\{}\begin{array}[]{ll}1&\quad\mbox{if $g_{r}=s$,}\\ 0&\quad\mbox{otherwise,}\\ \end{array}$ (20)
so the conventional mutual information of Eq. (13) is
$I_{0}(c;g)=\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}=H_{0}(g).$ (21)
This answer is not merely wrong; it is maximally so. The mutual information should take its minimum value of zero, but instead it takes the value $H_{0}(g)$, which is the maximum possible since $H_{0}(g)$ is an upper bound as we have said. The reason for this result is that in this case the contingency table itself uniquely defines $g$, so neglecting it puts the mutual information in error by an amount equal to the complete information cost of the ground truth. If we include the subleading terms on the other hand, this erroneous behavior disappears. Assuming for simplicity that the ground-truth groups are of equal size, the optimal concentration parameters are $\alpha_{g|c}=0$ and $\alpha_{g}=\infty$, and the reduced mutual information of Eq. (12) becomes
$=0.$ (22)
which is the correct answer.
@@ II.1 Reduced mutual information
A related shortcoming of the standard mutual information is that for finite $n$ a random labeling $c$ will have positive mutual information with respect to any ground truth in expectation: because mutual information is non-negative, fluctuations due to randomness in finite systems will produce non-negative values only and hence their average will in general be positive [4, 9]. This seems counterintuitive: we would expect the average value for a random labeling to be zero.
@@
We can rectify this issue by using another variant of the mutual information, which subtracts off the expected value of the measure, thereby making the average zero by definition. To do this we must first decide how a random labeling is to be defined—over what distribution of candidate labelings are we averaging? The conventional choice is to take the uniform distribution over labelings that share the same group sizes $n^{(c)}$ as the actual candidate $c$. This yields the adjusted mutual information of Vinh et al. [4]:
$I_{A}(c;g)=I_{0}(c;g)-\langle I_{0}(c;g)\rangle_{\{c|n^{(c)}\}},$ (23)
@@
where the expectation is over the relevant ensemble.
This measure can also be treated within the information-cost framework of the present paper, as described in [6]. There it is shown that the subtracted term $\langle I_{0}(c;g)\rangle_{\{c|n^{(c)}\}}$ is precisely equal to the average cost of transmitting the contingency table when labelings are drawn from the uniform distribution. Thus, one can view the adjusted mutual information as a reduced measure, but differing from the one proposed in this paper in two ways. First, the correction term counts the cost of transmitting the contingency table but not the cost of transmitting the counts $n^{(g)}$, as described in Section II. This, however, is only a minor concern—the latter is usually much smaller than the former and its neglect introduces only a small error in most cases.
@@
The second difference is more serious and revolves around the choice to draw labelings uniformly from those with given $n^{(c)}$. As discussed in [6], this distribution heavily favors contingency tables with relatively uniform entries, simply because there are many more labelings that correspond to uniform tables than to non-uniform ones. Real contingency tables, on the other hand, are often highly non-uniform, since many applications of the mutual information focus on labelings that are somewhat similar to the ground truth (producing a non-uniform table). In such cases, the average used in the adjusted mutual information puts most of its weight on configurations that are very different from those that occur in reality, making it a poor representation of true information costs. The measure proposed in this paper, by contrast, deliberately allows for non-uniform tables by drawing them from a Dirichlet-multinomial distribution and we argue that this is a strong reason to favor it over the adjusted mutual information. In addition it is also derived naturally from a fully information-theoretic argument, by contrast with the more ad hoc derivation of the adjusted mutual information. Notwithstanding, in Section III we give results using both reduced and adjusted mutual information, and find fairly similar outcomes in the two cases.
II.2 Normalization of the mutual information

### v3

@@ 1 Introduction
A common task in data analysis is the comparison of two different labelings of a set of objects. How well do demographics predict political affiliation? How accurately do blood tests predict clinical outcomes? These are questions about how similar one labeling of individuals/items/results is to another and they are commonly tackled using the information theoretic measure known as mutual information, in which an experimental or computational measure of some kind is compared against a ground truth.
@@
Mutual information works by asking of how efficiently we can describe one labeling if we already know the other [1]. Specifically, it measures how much less information it takes to communicate the first labeling if we know the second versus if we do not. As an example, mutual information is commonly used in network science to evaluate the performance of algorithms for network community detection [2]. One takes a network whose community structure is already known and applies a community detection algorithm to it to infer the communities. Then one uses mutual information to compare the output of the algorithm to the known correct communities. Algorithms that consistently achieve high mutual information scores are considered good. We will use this application as an illustrative example later in the paper.
@@
Mutual information has a number of appealing properties as a tool for comparing labelings. It is invariant under permutations of the labels, so that, for example, two divisions of a network into communities that differ only in the numbering of the communities will be correctly identified as being the same. It also returns sensible results in cases where the number of distinct label values is not the same in the two labelings. However, the standard mutual information measure also has some significant shortcomings, and two in particular that we highlight in this paper. First, it has a bias towards labelings with too many distinct label values. For instance, a community detection algorithm that incorrectly divides networks into significantly more groups than are present in the ground truth can nonetheless achieve a high mutual information score. A number of approaches for correcting this flaw have been proposed. One can apply direct penalties for incorrect numbers of groups [3] or subtract correction terms based on the average mutual information over some ensemble of candidate labelings [4, 5] or on the statistics of the contingency table [6, 7]. For reasons we discuss shortly, we favor the latter approach, which leads to the measure known as the reduced mutual information.
@@
The second drawback of the mutual information arises when the measure is normalized, as it commonly is to improve interpretability. The most popular normalization scheme creates a measure that runs between zero and one by dividing the mutual information by the arithmetic mean of the entropies of the two labelings being compared [8], although one can also normalize by the minimum, maximum, or geometric mean of the entropies. As we demonstrate in this paper, however, these normalizations introduce biases into the results by comparison with the unnormalized measure, because the normalization factor depends on the candidate labeling as well as the ground truth. This effect can be large enough to change scientific outcomes, and we provide examples of this phenomenon.
@@
In order to avoid this latter bias, while still retaining the interpretability of a normalized mutual information measure, we favor normalizing by the entropy of the ground-truth labeling alone. This removes the source of bias but introduces an asymmetry in the normalization. At first sight this asymmetry may seem undesirable, and previous authors have gone to some lengths to avoid it. Here, however, we argue that it is not only justified but actually desirable, for several reasons. First, many of the classification problems we consider are unaffected by the asymmetry, since they involve the comparison of one or more candidate labelings against a single, unchanging ground truth. Moreover, by contrast with the multitude of possible symmetric normalizations, the asymmetric measure we propose is the unique way to normalize the mutual information without introducing biases due to the normalization itself.
@@
More broadly, the asymmetric measure is more informative than the conventional symmetric one. Consider for instance the common situation where the groups or communities in one labeling are a subdivision of those in the other. We might for example label individuals by the country they live in on the one hand and by the town or city on the other. Then the more detailed labeling tells us everything there is to know about the coarser one, but the reverse is not true. Telling you the city fixes the country, but not vice versa. Thus one could argue that the mutual information between the two eshould be asymmetric, and this type of asymmetry will be a key feature of the measures we study.
@@
Both drawbacks of the standard mutual information—bias towards too many groups and dependence of the normalization on the candidate labeling—can be addressed simultaneously by using an asymmetric normalized reduced mutual information as defined in this paper. In support of this approach we present an extensive comparison of the performance of this and other variants of the mutual information in network community detection tasks, generating a large number of random test networks with known community structure and a variety of structural parameters, and then attempting to recover the communities using popular community detection algorithms. Within this framework, we find that conclusions about which algorithms perform best are significantly impacted by the choice of mutual information measure, and specifically that traditional measures erroneously favor algorithms that find too many communities, but our proposed measure does not. Code implementing our measure can be found at https://github.com/maxjerdee/clustering-mi.
2 Results
@@
Mutual information can be thought of in terms of the amount of information it takes to transmit a labeling from one person to another. We represent a labeling or division of $n$ objects into $q$ groups as a vector of $n$ integer elements, each with value in the range $1\ldots q$. We assume there to be a ground-truth labeling, which we denote $g$, with $q_{g}$ groups, and a candidate labeling for comparison $c$ with $q_{c}$ groups, generated for instance by some sort of algorithm. The mutual information $I(c;g)$ between the two is the amount of information saved when transmitting the truth $g$ if the receiver already knows the candidate $c$. We can write this information as the total entropy of $g$ minus the conditional entropy of $g$ given $c$:
I(c;g)=H(g)H(g|c).I(c;g)=H(g)-H(g|c). (1)
@@
Loosely, we say that $I(c;g)$ measures how much $c$ tells us about $g$. Since the conditional information content $H(g|c)$ is nonnegative, the mutual information must satisfy inequalities
$I(c;g)\leq H(g),H(c),$ (2)
which are saturated by
$I(g;g)=H(g),\qquad I(c;c)=H(c).$ (3)
Carefully accounting for all of the constituent information costs, we can compute the mutual information from [7]
@@
$=\log\frac{n!\prod_{rs}n_{rs}^{(gc)}!}{\prod_{r}n_{r}^{(c)}!\prod_{s}n_{s}^{(g)}!}+\log\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}-\sum_{r=1}^{q_{g}}\log\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1}$
$\qquad-\sum_{s=1}^{q_{c}}\log\binom{n_{s}^{(c)}+q_{g}\alpha_{g|c}-1}{q_{g}\alpha_{g|c}-1}+\sum_{r=1}^{q_{g}}\sum_{s=1}^{q_{c}}\log\binom{n_{rs}^{(gc)}+\alpha_{g|c}-1}{\alpha_{g|c}+1}$. (4)
In this expression the quantity $n_{r}^{(g)}$ denotes the number of objects in labeling $g$ that belong to group $r=1\ldots q_{g}$, and similarly for $n_{s}^{(c)}$. The quantities $n_{rs}^{(gc)}$ are the entries of the contingency table, and are equal to the number of objects that simultaneously belong to group $s$ in the candidate labeling $c$ and group $r$ in the ground truth $g$. Evaluating the sums and products over these table entries numerically, the value of $I(c;g)$ can be computed in time $\mathrm{O}(q_{c}q_{g})$. The parameter values $\alpha_{g}$ and $\alpha_{g|c}$ are chosen to optimize the transmission process using a fast line search, providing good compression for realistic labelings as discussed in [7].
The first term in Eq. (4) represents the information cost of transmitting the labelings given knowledge of the contingency table $n_{rs}^{(gc)}$. The remaining terms account for the information costs associated with the transmission of the table itself. In some cases the information required to transmit the $q_{g}\times q_{c}$ contingency table is significantly smaller than that needed for the $n$-vectors representing the labelings, and it is neglected in most treatments [6, 7], leading to the traditional mutual information measure, which we denote $I_{0}$:
@@
In other cases, however, the information cost of the contingency table can play a significant role and should be retained as in Eq. (4), which yields the quantity known as the reduced mutual information.
As an example, suppose a community detection algorithm simply places each node in its own group, resulting in a candidate labeling $c=(1,2,3,\ldots)$. No matter what the ground truth is, this choice of $c$ clearly contains no information about it whatsoever, so we might expect the mutual information to be zero. If we use the traditional mutual information, however, we get $I_{0}(c;g)=I_{0}(g;g)$, which is not merely nonzero, but implies that the candidate labeling is maximally informative, telling us everything about the ground truth. This is as wrong as it could be—in fact the candidate tells us nothing at all. The reduced mutual information, on the other hand, correctly says that $I(c;g)=0$ (see Section B in the Supplementary Materials).
For the same reason, the traditional mutual information gives a maximum score to any candidate $c$ which is a refinement of the true $g$, i.e., every community in $c$ is a subset of a community in $g$. More generally, the traditional mutual information is biased towards labelings with too many groups [3, 4, 6, 7, 9].
@@
A related shortcoming of the traditional mutual information is that for finite $n$ even a random labeling $c$ will have positive mutual information with respect to any ground truth in expectation: because the traditional mutual information is non-negative, fluctuations due to randomness will produce non-negative values only and hence their average will in general be positive [4, 10]. This seems counterintuitive; we would expect the average value for a random labeling to be zero.
@@
We can solve this problem by using another variant of the mutual information, which subtracts off the expected value, thereby making the average zero by definition. To do this we must first specify how the expectation is defined—over what distribution of candidate labelings are we averaging? The conventional choice is to take the uniform distribution over labelings that share the same group sizes $n^{(c)}$ as the actual candidate $c$. This yields the adjusted mutual information of Vinh et al. [4]:
$I_{A}(c;g)=I_{0}(c;g)-\bigl\langle I_{0}(c;g)\bigr\rangle_{\{c|n^{(c)}\}}\,$, (6)
@@
where the expectation $\langle\ldots\rangle$ is over the relevant ensemble.
@@
The adjusted mutual information can also be derived in a fully information-theoretic manner, as described in [6]. There it is shown that the subtracted term $\langle I_{0}(c;g)\rangle_{\{c|n^{(c)}\}}$ is precisely equal to the average cost of transmitting the contingency table when labelings are drawn from the uniform distribution. However, this distribution heavily favors contingency tables with relatively uniform entries, simply because there are many more labelings that correspond to uniform tables than to non-uniform ones. Real contingency tables, on the other hand, are often highly non-uniform, since applications of the mutual information focus on labelings that are somewhat similar to the ground truth (producing a non-uniform table). In such cases, the average used in the adjusted mutual information puts most of its weight on configurations that are very different from those that occur in reality, making it a poor representation of true information costs. The reduced mutual information considered in this paper, by contrast, deliberately allows for non-uniform tables by drawing them from a Dirichlet-multinomial distribution and we argue that this is a strong reason to favor it over the adjusted mutual information. Nonetheless, in Section 2.3 we give results using both reduced and adjusted mutual information, and find fairly similar outcomes in the two cases.
2.1 Normalization of the mutual information